if 4 is added to twice a number and this sum is multiplied by 2, the result is the same as if the number is multiplied by 3 and 4 is added to the product. what is the number?
Found 2 solutions by oberobic, ankor@dixie-net.com:
Answer by oberobic(2304)

2(2x+4) = 3x+4
4x + 8 = 3x + 4
x = -4
.
check this answer.
.
2(2(-4)+4) = 3(-4)+4
-16 +8 = -12 +4
-8 = -8
Correct.
.
Answer: x = -4
.
Done.

Answer by ankor@dixie-net.com(15660)

Let n = "the number"
:
Just write an equation for what it says.
"if 4 is added to twice a number and this sum is multiplied by 2, the result is
the same as if the number is multiplied by 3 and 4 is added to the product."
:
2(2n + 4) = 3n + 4
:
See if you can do this now.